OpenStudy (anonymous):

The equation x^4+px^3+qx^2+rx+s=0 has two roots which are reciprocols and the other two roots which are opposites. Show that r=ps.

ganeshie8 (ganeshie8):

say the roots are \(\alpha, \frac{1}{\alpha }, \beta , -\beta\)

ganeshie8 (ganeshie8):

\(x^4+px^3+qx^2+rx+s= (x-\alpha )(x-\frac{1}{\alpha })(x-\beta )(x + \beta)\)

ganeshie8 (ganeshie8):

multiply thru and compare \(x^3 , x \) coefficients and constant terms

OpenStudy (anonymous):

I did that but it doesn't seem to work out...Let me try get a picture of my working out.

ganeshie8 (ganeshie8):

you would get :- \(s = -\beta ^2\) \(p = -(\alpha + \frac{1}{\alpha })\) \(r = \beta^2 ( \alpha + \frac{1}{ \alpha } ) \)

ganeshie8 (ganeshie8):

so, r = ps. try again... post the picture il have a look :)

OpenStudy (anonymous):

sure, thanks a lot :D by the way, I'm new, how do you give 'medals'?

ganeshie8 (ganeshie8):

il message you about that... take a screenshot and npost the work... i cna have a quick look if u want... :)

ganeshie8 (ganeshie8):

looks good, maybe a small typo somewhere... let me check

ganeshie8 (ganeshie8):

got it, for 'x' coefficient, you need to take sum of products of 3 roots at a time

OpenStudy (anonymous):

but r=\[\alpha ^{2}\beta\] not \[\alpha ^{2}\beta+\frac{ \alpha ^{2} }{\beta}\]

OpenStudy (anonymous):

doesn't that cancel out to alpha^2*beta

ganeshie8 (ganeshie8):

if a, b, c, d are roots, then -r = abc + bcd + cda + dab

ganeshie8 (ganeshie8):

u need to take product of 3 roots at a time and take their sum, right ?

OpenStudy (anonymous):

OH i see what I did wrong I forgot one term hahah!! THANKS A LOT! :)

ganeshie8 (ganeshie8):

glad to hear :) nice work btw... yw !